# Analytics, have some humility: a statistical view of fourth-down decision making

### v4

@@ Appendix H Baseline coaches’ decision model
In Figure 14 we visualize the variable importance (via gain) of our ${\mathsf{XGBoost}}$ model. Interestingly, point spread has an extremely small impact on coaches’ fourth-down decisions. Perhaps this is because coaches don’t like to admit when their teams are underdogs as some sort of psychological leadership tool. We find, however, that point spread should impact fourth-down decision making. For instance, in certain game-states, it is advantageous for the favorites to be more aggressive (e.g., late in close games).
@@
Appendix I Evaluating coaches’ decision making
A football analyst shouldn’t penalize a coach for making a decision that has high uncertainty regardless of the effect size because the estimated edge may be due to noise. On this view, we should evaluate coaches only on plays for which we are confident (and perhaps also plays in which we lean towards a decision). In Figure 15 we rank coaches from 2018 to 2022 by the proportion of correctly made decisions that we are confident in. The top of this list has coaches like Kevin O’Connell and John Harbaugh from analytics bent organizations but also coaches that are considered more traditional on fourth down such as Ron Rivera, Sean McVay, and Mike Vrabel. While “analytics bent” coaches tend to follow fourth-down recommendations from win probability point estimates, which include recommendations on confident plays, some of the more traditional coaches seem to have a good feel for obvious fourth downs.
Figure 15: Ranking coaches by the proportion of fourth-down decisions they made correctly from 2018-2022 among fourth-down plays we are confident in.
Appendix J Additional example plays
Example play 3. In Figure 16 we visualize our decision procedure for a fourth-down play in which the Bears had the ball against the Jets in Week 12 of 2022. ${\mathsf{FG}}$ provides an edge over ${\mathsf{Go}}$ according to the ${\mathsf{WP}}$ point estimate ($+1.3\%$ ${\mathsf{WP}}$). But our $90\%$ confidence interval of the estimated gain in win probability by attempting a field goal is $[-3.2\%,3.9\%]$, indicating that ${\mathsf{FG}}$ could either be a great or a terrible decision. Also, more than half of our bootstrapped models say ${\mathsf{Go}}$ is better. In other words, we do not have enough data to be confident in our win probability point estimates, and we don’t know the optimal fourth-down decision at this game-state. Further, in the bottom right plot, notice how most of the colors are light. This indicates that the optimal decision is uncertain at most other combinations of yardline and yards to go at this game-state.
@@ Appendix I Evaluating coaches’ decision making
Figure 16: Our decision charts for example play 3. The bootstrapped models are split amongst ${\mathsf{Go}}$ and ${\mathsf{FG}}$ (${\mathsf{boot\%}}$, in the orange column in the top figure, is near $50\%$-$50\%$), so we don’t have enough data to be confident in the fourth-down recommendation provided by the point estimate (blue column in the top figure). Even though the pink dot in the lower left figure (representing the actual play’s yardline and yards to go) is in a yellow region (so the point estimate says ${\mathsf{FG}}$), the pink dot in the lower right figure is in a region of light color intensity (reflecting low ${\mathsf{boot\%}}$).
Example play 4. In Figure 17 we visualize our decision procedure for a fourth-down play in which the Ravens had the ball against the Jets in Week 1 of 2022. ${\mathsf{Punt}}$ provides a slight edge over ${\mathsf{Go}}$ according to the ${\mathsf{WP}}$ point estimate ($+1.2\%$ ${\mathsf{WP}}$). But $97\%$ of the bootstrapped models say ${\mathsf{Punt}}$ is better and our $90\%$ confidence interval of the estimated gain in win probability by punting is $[0.5\%,3.5\%]$, which is strictly positive. So, even if the edge is small, we are confident in this edge and recommend that the Ravens should ${\mathsf{Punt}}$. Further, in the bottom right plot, notice how most of the colors are dark outside of a large white boundary region. This indicates that we have higher certainty in our estimated optimal decision at this game-state.
@@
Figure 17: Our decision charts for example play 4. Nearly all the bootstrapped models say ${\mathsf{Punt}}$ is superior (${\mathsf{boot\%}}$, in the orange column in the top figure, is $97\%$), so we are confident that ${\mathsf{Punt}}$ has higher win probability than the other decisions. The pink dot in the lower left figure (representing the actual play’s yardline and yards to go) is in a red region (so the point estimate says ${\mathsf{Punt}}$) and the pink dot in the lower right figure is in a region of dark color intensity (reflecting high ${\mathsf{boot\%}}$).
Example play 5. In Figure 18 we visualize our decision procedure for an infamous fourth-down play in which the Raiders had the ball against the Rams in Week 14 of 2022. ${\mathsf{Go}}$ provides a strong edge over ${\mathsf{Punt}}$ according to the ${\mathsf{WP}}$ point estimate ($+3.9\%$ ${\mathsf{WP}}$). Further, $92\%$ of the bootstrapped models say ${\mathsf{Go}}$ is better and our $90\%$ confidence interval of the estimated gain in win probability by going for it is $[-0.1\%,4.4\%]$, which is nearly positive. Thus, we are confident in this edge, and we strongly recommend that the Raiders should ${\mathsf{Go}}$555 In real life, the Raiders punted. Then, Rams quarterback Baker Mayfield countered with a successful 98-yard drive to win the game.
@@
Figure 18: Our decision charts for example play 5. Nearly all the bootstrapped models say ${\mathsf{Go}}$ is superior (${\mathsf{boot\%}}$, in the orange column in the top figure, is $92\%$), so we are confident that ${\mathsf{Go}}$ has higher win probability than the other decisions. The pink dot in the lower left figure (representing the actual play’s yardline and yards to go) is in a green region (so the point estimate says ${\mathsf{Go}}$) and the pink dot in the lower right figure is in a region of dark color intensity (reflecting high ${\mathsf{boot\%}}$).
Appendix K Limitations and future work
Although our analysis improves the state of the art, it is not without limitations and there are many avenues for future work. To begin, even though the randomized cluster bootstrap produces substantially wide confidence intervals for win probability estimates, it underestimates uncertainty because it quantifies sampling uncertainty but not model uncertainty. The former is uncertainty in our point estimates resulting from fitting a model on a finite dataset (“variance”) and the latter is uncertainty caused by our model being wrong or biased (“bias”). In our simulation study from Section 3.1, there is no model uncertainty because we know true win probability is indeed a function of time, score differential, and field position. Win probability in real football, however, is highly likely a function of unobserved confounders. For example, how well Tom Brady slept the previous night could affect his team’s win probability. A more elaborate uncertainty quantification would capture model uncertainty.
In this work we used bootstrapping to drive uncertainty quantification. The way we tuned the bootstrap to achieve adequate marginal coverage in Section 3.1 is reminiscent of conformal inference. Conformal prediction intervals provide guarantees regarding the marginal coverage by ensuring every observation gets a holdout version of the interval (Lei et al.,, 2018). That the bootstrap methods with $\phi=1$ don’t achieve adequate coverage could be due to the time series dependence structure of football data. Recent work on conformal approaches for time series may help mitigate this issue (Angelopoulos et al.,, 2023). Nonetheless, valid distribution-free inference may not be possible in a binary outcome setting (Barber,, 2020). For instance, no confidence or prediction interval (e.g., $[0.63,0.71]$) will cover a 00 or $1$ outcome value.
The way we tuned the bootstrap is also reminiscent of random forest prediction intervals (Zhang et al.,, 2020). Random forest prediction intervals are based on the empirical distribution of out-of-bag prediction errors. These intervals can be obtained as a by-product of a single random forest and under regularity conditions have asymptotically correct coverage rates.
Our analysis captures uncertainty in first-down win probability but not in decision transition probabilities (conversion probability, field goal success probability, and expected next yardline after punting). These models are fit from thousands of play-level i.i.d. observations, and so are subject to little sampling uncertainty, but are subject to nontrivial model uncertainty, as they make simplifying assumptions. Conversion probability in particular is a delicate concept, as it depends on the offensive play call, the defensive play call, and the individual characteristics of each of the players on the field. A more fine-grained analysis would account for this additional uncertainty.
Finally, in this work we recommend a fourth-down decision when we are confident it has higher win probability than all other decisions. If teams follow our recommendations, their behavior will change and win probability will change accordingly. Statistical win probability models that learn from the game outcomes defining the recent history of football do not account for this distribution shift. State-space models, on the other hand, can account for these changes by altering the probability that a team goes for in on fourth down as a function of game-state. A more elaborate analysis would account for this distribution shift.

### v5

@@ Appendix D Baseline coaches’ decision model
In Figure 13 we visualize the variable importance (via gain) of our ${\mathsf{XGBoost}}$ model. Interestingly, point spread has an extremely small impact on coaches’ fourth-down decisions. We find, however, that point spread should impact fourth-down decision making. For instance, in certain game-states, it is advantageous for the favorites to be more aggressive (e.g., late in close games).
@@
Appendix E Additional example plays
Example play 3. In Figure 14 we visualize the decision procedure for a fourth-down play in which the Bears had the ball against the Jets in Week 12 of 2022. ${\mathsf{FG}}$ provides an edge over ${\mathsf{Go}}$ according to the ${\mathsf{WP}}$ point estimate ($+1.3\%$ ${\mathsf{WP}}$). But, the $90\%$ confidence interval of the estimated gain in win probability by attempting a field goal is $[-4\%,5\%]$, indicating that ${\mathsf{FG}}$ could either be a good or a bad decision. Also, about half of the bootstrapped models say ${\mathsf{Go}}$ is better. In other words, we do not have enough data to be confident in the win probability point estimates, and we don’t know the optimal fourth-down decision at this game-state. Further, in the bottom right plot, notice how most of the colors are light. This indicates that the optimal decision is uncertain at most other combinations of yardline and yards to go at this game-state.
@@ Appendix E Additional example plays
Figure 14: Decision charts for example play 3.
Example play 4. In Figure 15 we visualize the decision procedure for a fourth-down play in which the Ravens had the ball against the Jets in Week 1 of 2022. ${\mathsf{Punt}}$ provides a slight edge over ${\mathsf{Go}}$ according to the ${\mathsf{WP}}$ point estimate ($+0.08\%$ ${\mathsf{WP}}$). But most of the bootstrapped models say ${\mathsf{Punt}}$ is better and the $90\%$ confidence interval of the estimated gain in win probability by punting is $[-1\%,3\%]$, which is largely positive. So, even if the edge is small, we are confident in this edge and recommend that the Ravens should ${\mathsf{Punt}}$. Further, in the bottom right plot, notice how most of the colors are dark outside of a large white boundary region. This indicates that we have higher certainty in the estimated optimal decision at this game-state.
@@
Figure 15: Decision charts for example play 4.
Example play 5. In Figure 16 we visualize the decision procedure for an infamous fourth-down play in which the Raiders had the ball against the Rams in Week 14 of 2022. ${\mathsf{Go}}$ provides a strong edge over ${\mathsf{Punt}}$ according to the ${\mathsf{WP}}$ point estimate ($+4.1\%$ ${\mathsf{WP}}$). Further, many of the bootstrapped models say ${\mathsf{Go}}$ is better and the $90\%$ confidence interval of the estimated gain in win probability by going for it is $[-1\%,5\%]$, which is largely positive. Thus, we recommend that the Raiders should ${\mathsf{Go}}$666 In real life, the Raiders punted. Then, Rams quarterback Baker Mayfield countered with a successful 98-yard drive to win the game.
@@
Figure 16: Decision charts for example play 5.
Appendix F The difficulty of formulating a probabilistic state-space win probability model
Proprietors of probabilistic models believe that introducing bias in order to reduce variance improves the overall accuracy of the resulting win probability estimator. Nevertheless, these models are subject to their own set of issues, and we believe they aren’t as low-variance as some analysts claim. Properly modeling the distribution of the outcome of a play is nontrivial. In contrast to the simple binary win/loss outcome of statistical win probability models, the outcome variable of a transition probability model is the next game-state, which could include a change in yardline, score, time, or timeouts. This distribution is quite complex: there is a spike at gaining 0 yards for incompletions, a spike for a touchdown, spikes for penalties, and other smooth possibly multimodal distributions for the outcome of run or pass plays, each of which change as a function of team quality and other confounders Biro and Walker, 2023a ; Biro and Walker, 2023b . Typical transition probability models are riddled with selection bias, as a coach generally chooses play calls that work for his specific players and don’t generalize to an “average” team. Further, uncertainty in these transition probabilities may, after being properly propagated through a state-space model, result in just as much (if not more) uncertainty in estimated win probability than estimates from statistical models. Additionally, one must take great care to carefully encode all the subtle rules of football into her model, and one needs sufficient computing power to simulate enough games to estimate win probability with enough granularity. We look forward to a public facing exploration of probabilistic win probability models in the future.